what is the equation of 7f+18=4f-33
Answer by pwac(253)

7f+18=4f-33
what you have here is already an equation.
I think you possibly want the value of f
First subtract 4f from both sides
7f+18-4f=4f-33-4f
3f+18=-33 now subtract 18 from both sides
3f+18-18=-33-18
3f=-51 divide both sides by 3 to leave:
f=-17
Pete